# Jailbreak Attacks and Defenses against Multimodal Generative Models: A Survey

### v1

@@ V-C Evaluation Metric
V-C1 Metrics for Jailbreak Attack
Evaluating jailbreak attacks on multimodal models requires metrics like Attack Success Rate, which quantifies the proportion of successful attacks that bypass the model’s safety mechanisms.
$\bullet$ Attack Success Rate. The Attack Success Rate (ASR) is a commonly used metric for evaluating the effectiveness of jailbreak methods. It quantifies the proportion of successful attacks out of the total number of attempts. Formally, let $N_{total}$ denotes the total number of jailbreak inputs and $N_{success}$ denotes the number of successful attacks. The ASR can then be defined as:
@@ V-C1 Metrics for Jailbreak Attack
$\bullet$ Prompt Perplexity. Perplexity (PPL) [160] serves as a metric for evaluating the readability and linguistic fluency of jailbreak prompts. Many adversarial prompts targeting MLLMs tend to contain garbled or nonsensical characters, which makes them easily detectable and filtered by defense methods that rely on high-perplexity detection. As a result, attack methods that generate low-perplexity prompts are becoming increasingly noteworthy. Formally, let $W=\left(w_{1},w_{2},...,w_{n}\right)$ denote a text sequence where $w_{i}$ represents the $i$-th token. The perplexity of the sequence $W$ is defined as:
@@
V-C2 Metrics for Jailbreak Defense
Evaluating jailbreak defenses in MLLMs often involves monitoring changes in the Attack Success Rate. By comparing ASR before and after the implementation of defense mechanisms against jailbreak prompts, researchers can effectively assess the defenses’ impact. A substantial reduction in ASR suggests that the defense effectively reduces the model’s vulnerability to adversarial inputs, demonstrating its robustness in preventing harmful or unsafe outputs.
For T2I models, Frechet Inception Distance and CLIP Score are utilized to evaluate both the visual quality of generated images and their semantic alignment with adversarial intent following the implementation of defense mechanisms.
@@ V-C2 Metrics for Jailbreak Defense
$\bullet$ CLIP Score. CLIP Score [162] measures the alignment of the generated images with their corresponding descriptions. This approach first utilizes the CLIP model $E_{clip}$ to obtain the semantic embeddings of adversarial images. The CLIP Score is then calculated by computing the feature similarity between malicious adversarial images $\mathcal{Y}_{adv}$ and the clean prompts $\mathcal{X}_{clean}$ modified from the toxic content:
$CLIP\,\,Score=Cos\left(E_{clip}\left(\mathcal{Y}_{adv}\right),E_{clip}\left(\mathcal{X}_{clean}\right)\right)$ (14)
@@ VI Future Work
As previously discussed, various studies have investigated jailbreak attacks and defenses for multimodal generative models. However, this area remains in its early stages, with many challenges and potential vulnerabilities yet to be fully explored. This section presents an in-depth analysis of the open challenges and future research directions in multimodal model jailbreaks, aiming to drive further progress toward developing more secure generative systems.
@@ VI-A Multimodal Jailbreak Attack
VI-A1 Underexplored Video/Audio Vulnerabilities
While significant progress has been made in researching jailbreak attacks and defenses for text- and image-based models, video and audio modalities remain underexplored. These modalities are crucial for real-world multi-modal communication, yet their vulnerabilities in generative models, particularly for understanding and generation, are not well understood. The complex temporal dynamics and acoustic features inherent to video and audio pose distinct challenges for evaluating and addressing jailbreak attacks, necessitating more focused research.
VI-A2 Interleaving Multimodal Input
Interleaving multimodal jailbreak attacks pose a significant challenge, as they can simultaneously exploit content across multiple modalities, such as text, images, and audio. While existing research on multimodal jailbreak attacks often involves input from multiple modalities, the attack strategies primarily exploit vulnerabilities in one specific modality, such as text [79, 69, 76, 75] or image [18, 83, 86, 71, 72], rather than addressing the complexities of cross-modal interactions. Therefore, exploring robust techniques for handling interleaving multimodal jailbreak inputs is a crucial future research direction. Conducting interleaving jailbreak attacks across multiple modalities includes the interactions between modalities to simultaneously exploit vulnerabilities in a unified manner. The complexity arises not only from managing multiple data types but also from synchronizing them to achieve a cohesive malicious outcome. This collaborative nature makes detection and defense significantly more difficult, as traditional models are often optimized to handle single-modal inputs.
VI-A3 Diverse Multimodal Output
Attacks targeting Any-to-Any models remain largely unexplored, yet the potential for jailbreak attacks on these models is significant, as they can generate multimodal outputs simultaneously. For instance, a jailbreak attack might involve generating instructions on how to create a bomb, not only producing a textual description, as seen in existing any-to-text models [79, 69, 76, 75, 71, 72], but also generating multiple corresponding images or even sequential content resembling a comic strip. This capability greatly increases the harmful potential of such attacks, as it allows the adversary to convey complex and dangerous information across multiple modalities.
@@ VI-B Multimodal Jailbreak Defense
VI-B1 Developing a Multi-strategy Defense Mechanism
Defense methods based solely on fine-tuning cannot cover a wide range of threats while training-free methods often lack effectiveness and accuracy. Discriminative approaches are also slow to adapt to the evolving strategies of attackers and are therefore incomplete. Future defense strategies can integrate multimodal information to build a robust, layered defense system. This system would combine techniques such as prompt revision, model fine-tuning, and generation guidance to create a more comprehensive and adaptable defense mechanism.
VI-B2 Defense Cost vs. Performance Trade off
Real-time detection of input and generated content introduces significant computational overhead, especially in complex multimodal systems where handling cross-modal information affects inference efficiency. Enhancing defense robustness often comes with high computational costs, potentially degrading model performance. Future research should focus on designing efficient and lightweight defense solutions that maintain strong security without compromising system performance.
VI-B3 Transparency and Explainability of Defense Mechanisms
As multimodal defense mechanisms become more complex, ensuring their transparency and explainability will be a key research focus. Future work should explore ways to make defense systems more transparent, enabling users and developers to understand how these mechanisms function, along with their limitations. Enhancing transparency will improve the credibility and trustworthiness of defense methods.
VI-B4 Research on Personalized Defense Schemes
To address the varying security needs of different users and application scenarios, future research can focus on developing personalized defense schemes. These schemes would allow the strength of defense policies to be dynamically adjusted based on user preferences, specific use cases, and content security levels. For example, stricter defense mechanisms could be activated for sensitive content, while maintaining flexibility and variety for general users.
@@ VI-C Multimodal Jailbreak Evaluation
VI-C1 Limitation in diverse harmful concepts
A key challenge in the evaluation of multimodal models, particularly for T2I systems, is the limited scope of current specialized detectors, which primarily focus on two harmful concepts: violence and pornography. This narrow focus restricts the automated assessment of a broader range of harmful content, such as hate speech, misinformation, and cybercrime. Future work should aim to expand the range of automated detectors to encompass a more comprehensive array of harmful categories. By developing detectors that cover diverse harmful concepts, we can ensure more robust safety evaluations, addressing the complex challenges posed by multimodal models in real-world applications.
VI-C2 Evaluation criteria of multimodal security
Currently, there is no unified evaluation standard in the field of multimodal defense, and the effectiveness of various defense methods cannot be measured systematically. Future work can focus on building a more sophisticated multimodal security evaluation framework to evaluate how defense methods actually perform under different attack scenarios.

### v2

@@ V-C Evaluation Metric
Evaluation metrics can be broadly categorized into two primary types: robustness metrics and utility metrics. Robustness metrics are designed to assess the resilience of a model against jailbreak inputs. Utility metrics evaluate the relevance, accuracy, and overall quality of the content generated by the model.
C.1 Robustness Metrics
The attack success rate (ASR) is a widely used metric to evaluate the robustness of attack and defense. In the context of the attack, a higher ASR indicates that the generated content successfully bypasses model safeguards, leading to harmful outputs. On the defense side, a reduction in ASR suggests that the defense mechanisms are effective in preventing the generation of unsafe content.
$\bullet$ Attack Success Rate. The Attack Success Rate (ASR) quantifies the proportion of successful attacks out of the total number of attempts. Formally, let $N_{total}$ denotes the total number of jailbreak inputs and $N_{success}$ denotes the number of successful attacks. The ASR can then be defined as:
@@ C.1 Robustness Metrics
C.2 Utility Metrics
Utility Metrics include indicators designed to evaluate content quality across different modalities. Specifically, Prompt Perplexity measures the coherence and fluency of textual prompts. For visual content, Frechet Inception Distance measures the visual quality of generated images, and the CLIP Score evaluates the semantic alignment between images and their target intent.
$\bullet$ Prompt Perplexity. Perplexity (PPL) [174] serves as a metric for evaluating the readability and linguistic fluency of jailbreak prompts. Many adversarial prompts targeting MLLMs tend to contain garbled or nonsensical characters, which makes them easily detectable and filtered by defense methods that rely on high-perplexity detection. As a result, attack methods [70, 102, 66] that generate low-perplexity prompts are becoming increasingly noteworthy. Formally, let $W=\left(w_{1},w_{2},...,w_{n}\right)$ denote a text sequence where $w_{i}$ represents the $i$-th token. The perplexity of the sequence $W$ is defined as:
@@ C.2 Utility Metrics
$\bullet$ CLIP Score. CLIP Score [176] measures the alignment of the generated images with their corresponding descriptions. This approach first utilizes the CLIP model $E_{clip}$ to extract the semantic embeddings of adversarial images. The CLIP Score is then obtained by calculating the cosine similarity between malicious adversarial images $y_{adv}$ and the clean prompts $x_{clean}$ modified from the toxic content:
$CLIP\,\,Score=Cos\left(E_{clip}\left(y_{adv}\right),E_{clip}\left(x_{clean}\right)\right)$ (17)
@@ VI Future Work
Jailbreak attack and defense against multimodal generative models remains a very challenging and open research task. In this section, we share key insights into future research directions for multimodal model jailbreaks, pinpointing what is missing in the current research and identifying directions worth further exploration.
@@ VI-A Multimodal Jailbreak Attack
$\bullet$ Expand Multimodal Vulnerabilities. With the advancement of multimodal generative models, input-output configurations have expanded to include video and audio modalities, which remain relatively underexplored in the context of jailbreak attacks and defenses. Unlike text and images, the temporal dependencies of video frames and the inherent acoustic features of audio present unique challenges in both launching and defending against jailbreak attacks. There is an urgent need for further research into the vulnerabilities of models in the video and audio modalities, which will facilitate the development of more robust safety alignment mechanisms.
$\bullet$ Joint Modal Optimization. Another interesting research direction is to explore the joint optimization between multiple modalities to achieve competitive attack performance more deceptively. Several pioneering studies have been conducted, e.g., [95] executes jailbreaks by jointly optimizing textual and visual prompts. Under this context, the adversary does not treat each modality in isolation but instead looks for strategies in which the different modalities can interact to exploit model vulnerabilities in a unified manner.
$\bullet$ Combined Multimodal Output. Most existing research focuses on either any-to-text [79, 62, 63, 64, 80, 74, 71, 55, 72, 73, 18] or any-to-vision models [66, 67, 82, 69, 85, 70, 52, 76, 77, 78, 19], primarily addressing single-modality output configurations. However, the potential for harmful multimodal outputs generated by any-to-any models remains largely unexplored. Compared with single-modality jailbreak outputs, combined text-image-video-audio outputs can amplify the credibility and psychological influence of the harmful content. For instance, malicious textual descriptions paired with manipulated visual content, synchronized audio commentary, and supporting video sequences, can collaboratively create a unified and deceptive message.
@@ VI-B Multimodal Jailbreak Defense
$\bullet$ Hybrid Defense System. Defense methods relying solely on fine-tuning models often fail to adapt to the evolving strategies of attackers. Conversely, training-free methods frequently fall short in delivering the necessary effectiveness and accuracy. Future research should prioritize the development of hybrid defense systems that integrate diverse techniques to enhance robustness and adaptability. Such a system would combine external refiners with internal tuning-based and guidance-driven strategies to promote collaboration.
$\bullet$ Overall Quality Preserving. Fine-tuning pre-trained generative models to erasure unsafe concepts has shown promising progress in defending against jailbreak attacks. However, the process of erasing or unlearning a large number of concepts may result in a degradation of overall generation quality [121, 133, 142, 130] including reduced coherence, fidelity and diversity in generated outputs. To mitigate these challenges, there is growing interest in developing adaptive forgetting mechanisms that selectively target unsafe concepts while preserving adjacent knowledge, as well as quality restoration strategies to maintain the quality of generative outputs.
$\bullet$ Transparency and Explainability. As defense mechanisms become increasingly complex, ensuring their transparency and explainability has become a critical research focus. Some pioneer efforts such as GuardT2I [106], have explored optimizing language models to interpret the malicious intent behind adversarial inputs. However, further endeavors are imperative to enable both users and developers to comprehensively understand the inner workings of these defense mechanisms, including their strengths and limitations.
$\bullet$ Personalized Defense Schemes. To address the varying security needs of different application scenarios, future research should focus on developing personalized defense schemes. First, enhanced stringency for sensitive content in high-stakes scenarios such as financial systems, healthcare applications, or government communications. Second, balancing security with the need for broader generative variability in scenarios that prioritize creative outputs, such as content generation for entertainment or education. These schemes would allow the security levels of defense policies to be dynamically adjusted.
@@ VI-C Multimodal Jailbreak Evaluation
$\bullet$ Scope of Specialized Detectors. Current specialized detectors for automatic evaluation primarily focus on a limited set of harmful concepts, such as violence and pornography. This narrow scope significantly limits the ability to assess a broader spectrum of harmful content, including hate speech, misinformation, and cybercrime. Future research should prioritize expanding the capabilities of automated detectors to cover a more comprehensive range of harmful categories. By developing detectors capable of addressing diverse harmful concepts, the safety evaluation of multimodal models can become more robust and better aligned with the complex challenges encountered in real-world applications.
$\bullet$ Comprehensive Evaluation Benchmark. The robustness of multimodal defense mechanisms is typically evaluated using a limited set of attack methods, which often falls short of comprehensively assessing their effectiveness [138, 133, 131, 128, 61]. To address this, future research should prioritize the development of a comprehensive benchmark dataset that incorporates a wide variety of attack methods across diverse scenarios. The establishment of a standardized evaluation framework would not only promote consistency in research methodologies but also facilitate fair comparisons between defense strategies.